OpenStudy (anonymous):

adam has dimes and quarters in his pocket. there are 28 coins in all, and they are worth $4.60 altogether. how many quarters does he have?

OpenStudy (anonymous):

let d and q = the number of units. d + q = 28 0.10d + 0.25q = 4.60 => the values of 0.10 and 0.25 are the values of a dime and a quarter

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

you only need to find q, so we can work on that one first. d = 28 - q [from the first equation] plug that d in the second equation: 0.10(28 - q) + 0.25q = 4.6 the equation is now only in terms of q and you can solve. that's it. if you want to find how many dimes, you can plug q in the first equation. lmk if i should keep solving or if you got it :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

glad i could help :)